The Internet Book Company charges $6 postage for 2 books, and $10 postage for 5 books.
a) Find an equation relating the postage charge and number of books.
b) How much will it charge to post 8 books?

Answer by John10(297)

The Internet Book Company charges $6 postage for 2 books, and $10 postage for 5 books.
a) Find an equation relating the postage charge and number of books.
b) How much will it charge to post 8 books?
-------------------------------------------------------------------
Let x be the number of books and y be the cost of books
a.Find an equation relating the postage charge and number of books.
We have two points (2, 6) and (5, 10)
The slope is:
m = (10 - 6)/(5 - 2) = 4/3
The equation will be:
y - 6 = 4/3(x - 2)
y - 6 = 4x/3 - 8/3
y = 4x/3 + 6 - 8/3
y = 4x/3 + 10/3
b.How much will it charge to post 8 books?
We know the equation
y = 4x/3 + 10/3
If x = 8 then
y = 4(8)/3 + 10/3 = 32/3 + 10/3 = $14
So it will charges $14 for 8 books.
John10